Determine whether the given pairs of lines are parallel, coincident or intersecting?
L: x+y=10
M: 3x+3y=6

Answer by stanbon(75887)

Determine whether the given pairs of lines are parallel, coincident or intersecting?
L: x+y=10
M: 3x+3y=6
-----
Put each in the slope-intercept form:
y = -x + 10
y = -x + 2
-----
Both have slope = -1 and different y-intercepts.
---
So the lines are parallel.
==================
Cheers,
Stan H.